​Campus Stop, Inc. is a student co-op. Campus Stop uses a perpetual inventory system


Campus Stop, Inc. is a student co-op. Campus Stop uses a perpetual inventory system. The following transactions (summarized) have been selected for analysis 

a. Sold merchandise for cash (cost of merchandise $147.990).  $266,200 

b. Received merchandise returned by customers an unsatisfactory (but in perfect condition) for cash refund (original cost of merchandise $740).   1,690

c. Sold merchandise (costing $7,200) to a customer on account with term n/30.   16,000

d. Collected half of the balance owed by the customer in (c).                               8,000

 e. Granted a partial allowance relating to credit sales the customer in (e) had not yet paid.          1,720


Required:

 1. Compute Net Sales and Gross Profit for Campus Stop
Answer #1
1
Gross sales 282200 =266200+16000
Less: Sales return and allowance 3410 =1690+1720
Net sales 278790
Net sales 278790
Less: Cost of goods sold 154450 =147990-740+7200
Gross Profit 124340